OpenStudy (atlchic):

x=y and x+y=7, solve by elimination.

OpenStudy (anonymous):

Ah, elimination, ok

OpenStudy (atlchic):

we doing elimination

OpenStudy (anonymous):

Then, subtract y from both sides in the first equation: x-y = 0. Add both equations: 2x + 0y = 7 x = 7.2

OpenStudy (anonymous):

7/2*

OpenStudy (atlchic):

7/2+y=0, solve

OpenStudy (anonymous):

That means that y = -7/2. But substituting back, we would get: 7/2 - y = 0 => y = 7/2

OpenStudy (atlchic):

i need to find y

OpenStudy (anonymous):

It's a similar process: subtract y on both sides of equation 1 x-y = 0 Multiply by -1 -x + y = 0 Add both equations: 0x + 2y = 7 y = 7/2

OpenStudy (atlchic):

solve belimination, 2x-y=3 and 2/3x-y=-1

OpenStudy (atlchic):

by elimination